OpenStudy (anonymous):

4/5 = x/75 pls help me im desperate

OpenStudy (anonymous):

Multiply both sides by 75.

OpenStudy (anonymous):

k so do i do anything after tht?

OpenStudy (anonymous):

Just try to simplify the answer.

OpenStudy (anonymous):

k wait um i just learned this so when u say both sides?

OpenStudy (anonymous):

After you multiply by 75, you get: \[{4 \over 5}(75)=x \]

OpenStudy (anonymous):

so um uh wow if i had any idea it was this hard id ask my teacher so x= 75?

OpenStudy (anonymous):

x=75 times 4/5

OpenStudy (anonymous):

ohhhhhh k thank u sooo much i hate my math hw this is just the 2nd prob so im in 4 it lol

OpenStudy (anonymous):

You should simplify it actually, your answer should be x=(75/5)(4)=15*4 --> x=60. So your final answer is x=60

OpenStudy (anonymous):

so i got my anwnser and i think x=300

OpenStudy (anonymous):

300 is too much :)

OpenStudy (anonymous):

oh so even if its simplifiying prportions

OpenStudy (anonymous):

Which grade are you in?

OpenStudy (anonymous):

7th

OpenStudy (anonymous):

AnwarA u there?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

You still have another question?

OpenStudy (anonymous):

yah but its alot and i dont wanna bother u

OpenStudy (anonymous):

It's alright. I still have got few minutes.

OpenStudy (anonymous):

really thanks!

OpenStudy (anonymous):

7/8 = 56/n a. 8n=56 b. 7n=56 c. 7n=448 or d. none of these

OpenStudy (anonymous):

Have you heard about cross multiplication?

OpenStudy (anonymous):

yah but im not 2 good @ it

OpenStudy (anonymous):

It's ok. You soon will be.

OpenStudy (anonymous):

Let me rewrite it: \[{7 \over 8}={56 \over n}\] Now multiply the top of the left side by the bottom of the right side, and the top of the right side by the bottom of the left side.

OpenStudy (anonymous):

i did 58\[\times\] 8 and gt 464

OpenStudy (anonymous):

So you multiply 7 by n and 56 by 8, you then get: 7n=56*8 But as you said 56*8=448. So the answer would be: 7n=448

OpenStudy (anonymous):

Is that clear?

OpenStudy (anonymous):

ohhh thank u

OpenStudy (anonymous):

You're welcome. Wish you the best luck.

OpenStudy (anonymous):

thanks

OpenStudy (gina):

\[4/5=x/75, 5x=4*75 then x =4*75/5\]

OpenStudy (gina):

sportsbabe get your answer